OpenStudy (anonymous):

Evaluate each expression for m = 8 and n = 2. a. 2m - 3n b. 12 divided by n + 12m c. mn(to the power of 2)

OpenStudy (anonymous):

\[2\times 8-3\times 2\]for the first one

OpenStudy (anonymous):

second one is not clear

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

\[\frac{12}{2}+12\times 8\] maybe but not clear exactly what the expression is use the equation tool

OpenStudy (anonymous):

\[mn^2=8\times 2^2\]

OpenStudy (anonymous):

\[12 \div n + 12m\]

OpenStudy (triciaal):



OpenStudy (anonymous):

oops \[12\div 2+12\times 8\]